a glass vessel whose bottom has holes with diameter 0.1mm is filled with water what is the maximum height upto which water can be filled without leakage given surface tension=75*10^-3N/m and g=10m/s^2


 

Grade:12

4 Answers

Chetan Mandayam Nayakar
11 years ago

let pho be density of water. Let maximum height=h,let d=hole diameter

rho*gh*pi*d2/4=pi*d

Akash Kumar Dutta
11 years ago

apply
total pressure=pressure due to surface tension,

Chetan Mandayam Nayakar
11 years ago

my previous solution had a mistake:

rho*g*h*pi*d2/4=pi*d*surface tension

harshini k
11 years ago

soeery for the wrong title it must be sueface tension instead of fluid dynamics